OpenStudy (anonymous):

Evaluate the expression. (-3^2)^2 + (4-15)^3 * 5^-2

OpenStudy (anonymous):

=( -9)^2 + (-11)^3* 1/25 =81 + (-1331 * 1/25) =81 - 53.24 =27.76 or 3356/25

OpenStudy (anonymous):

Thank you @Caspia

OpenStudy (anonymous):

Welcome @StudiousStudents